# Vista: A Generalizable Driving World Model with High Fidelity and Versatile Controllability

### v1

@@ 3.1 Phase One: Learning High-Fidelity Future Prediction
Dynamics Enhancement Loss. Unlike general videos that cover rather small spaces, driving videos encompass much larger scenes. In most driving videos, distant and monotonous regions dominate the view, with the moving instances in the foreground only occupying a relatively small area. However, the latter often exhibit higher stochasticity, complicating their prediction. Since Eq. 1 supervises all outputs uniformly, it cannot effectively discriminate the nuances of different regions as Fig. 4(b) shows. As a result, the model cannot efficiently learn to predict realistic dynamics in crucial regions.
@@ Appendix A Discussions
(1) As a generalizable predictive model, Vista could be utilized as a forward dynamics model [13, 23] to simulate short-term dynamics and assist long-horizon planning tasks like visual navigation [99]. (2) It is also intriguing to utilize Vista as an implicit driving policy, which is spontaneously acquired through future prediction [1, 22]. After synthesizing the video plan, we can convert the resultant image trajectory to executable actions by a non-causal inverse dynamics model [4, 65], which can be efficiently learned from suboptimal data in contrast to the imitation learning pipeline [2, 127]. In autonomous driving, the inverse dynamics model could be implemented with visual odometry [69]. (3) In collaboration with the reward function, it is also worth investigating if Vista could facilitate model-based reinforcement learning by boosting the sampling efficiency in real-world scenarios [131].
@@
These two works have fundamental differences in control versatility and prediction fidelity. First of all, Vista is a generalizable world model that can be controlled by multi-modal action conditions. Although GenAD has also trained a trajectory-conditioned extension, its weights are fully fine-tuned on nuScenes and the generalization of its action control has never been validated. In contrast, Vista integrates versatile action controllability that can generalize to new scenarios in a zero-shot manner. Unlike GenAD that requires labeling OpenDV-YouTube with commands and texts, our collaborative training strategy skillfully averts this labor that may incur accumulated noises and conflicts [130]. In addition, Vista (10 Hz, 576$\times$1024) operates at much higher frame rate and resolution, considerably beyond the capability of GenAD (2 Hz, 256$\times$448) in both temporal and spatial axes. Different from GenAD, we also put forth several dedicated designs for high-fidelity prediction. We find that Vista, with a lower model complexity, achieves much better FID and FVD scores than GenAD (see Table 2).
@@
As one of the pioneering efforts, Vista still has a few limitations that call for future works. (1) Since Vista predicts futures at an exceptional spatiotemporal resolution, it is inevitable to be computationally expensive, particularly in downstream applications. Potential solutions may include faster sampling methods [83, 139] and training-based distillations [85, 87, 95, 97, 103, 117]. (2) It is possible that the prediction may undergo an apparent degradation during long-horizon rollouts or drastic view shifts. Extra refinements on the prediction results [6, 45, 50, 93, 119, 138] could be helpful. Speculatively, applying our recipe to more scalable architecture [50, 91] is also promising to address this limitation. (3) Similar to other controllable video generation methods [121], the chance of failure still persists in our action controls, especially for ambiguous intentions such as commands and goal points as Fig. 8 reveals. Incorporating more datasets with action annotations [11, 105, 122] for collaborative training could be beneficial. Using compositional classifier-free guidance [8, 18, 29] to amplify the individual impact of action conditions may also help (at a cost of increased inference compute). (4) Although our training data is based on the largest public driving dataset [130], it is nowhere near the entirety of Internet driving data, thus leaving huge untapped potential to further expand the capabilities of Vista.
@@ B.1 World Models
Intelligent agents should be able to make efficacious decisions even under unseen situations [9, 53, 70, 106, 132, 142]. This requires fundamental knowledge of the world that generalizes to rare cases. As an internal manifestation of such knowledge, a world model predicts plausible futures of the world given potential actions [9, 37, 70, 90, 106, 131]. In principle, it not only predicts how the environment will unfold over time, but also deduces the underlying physical dynamics and agentic behaviors. Such properties can be useful for representation learning [32, 41, 82, 98, 124], model-based reinforcement learning [36, 37, 39, 40, 88, 90, 92], as well as model-predictive control [24, 27, 32, 38, 86, 136]. Recent works [34, 74, 79, 80] also induce language-based world models from large language models, but are restricted in textual space and struggle with grounding on physics [23, 55].
Although world models have been extensively applied and made significant revolutions in simulated games [37, 39, 40] and indoor embodiment [66, 86, 113], such investigations for autonomous driving still lag behind [120, 137]. Different from other tasks, world modeling for autonomous driving poses unique challenges, which primarily arise from the large field of views with highly dynamic motions. Some practices imagine the world in the bird’s eye view (BEV) space [28, 48, 49, 72, 75]. Recent practices model the world state as raw sensor observations such as point clouds [7, 62, 134, 137, 140] and images [50, 57, 63, 84, 96, 118, 120, 124]. The latter category, namely visual world models, hold more promise for scaling up due to sensor flexibility and data accessibility. Nevertheless, existing methods are still restricted to a particular dataset [57, 71, 84, 118, 120, 137, 140] or simulator [7, 124], compromising their generalization ability to novel domains. Meanwhile, these efforts lack systematic designs for the driving domain and only model the world at relatively low frame rates and resolutions, which discards the fine-grained details and impairs their ability to express real-world behaviors. Moreover, most of them are restricted to a specific control modality [50, 57, 84, 118], which hinders the accommodation to prevailing planning algorithms [12, 14, 19, 52, 54, 60] and extension to more applications like decision-making [120] or user interaction [68]. Besides, existing methods seldom explore zero-shot action controllability across different datasets. The inferior generalization, fidelity and controllability collectively preclude existing driving world models from broadly facilitating the development of autonomous driving.
@@ C.2 Dataset
We utilize a rigorously filtered set of OpenDV-YouTube [130] for training, and incorporate nuScenes training set [10] during the action control learning phase. Concretely, we manually eliminate 15 hours of irrelevant content from OpenDV-YouTube, yielding approximately 1735 hours of unlabeled driving videos. Since nuScenes is heavily biased [77, 100, 120], we balance its samples based on command categories to foster the learning of rare actions. The video clips are sampled with 25 frames at 10 Hz. Although nuScenes [10] is logged at 12 Hz, we find no negative impact of treating them as 10 Hz videos. The model inputs are composed by cropping and resizing these clips to the target resolution. To categorize actions into commands, we follow the established conventions in planning [52, 54, 60] and define the command of ego-vehicle as "turn right" or "turn left" when its displacement exceeds 2 meters in the orthogonal direction relative to its initial heading. To allow more precise categorization, we additionally introduce a "stop" command when the forward driving distance is less than 2 meters.
@@ D.2 Additional Ablation Studies
Table 4: Effect of action independence. Without lossing generality, we choose the trajectory as a representative action for evaluation. The proposed constraint expedites the learning of actions.

### v2

@@ 3.1 Phase One: Learning High-Fidelity Future Prediction
Dynamics Enhancement Loss. Unlike general videos that cover rather small spaces, driving videos capture much larger scenes [133]. In most driving videos, distant and monotonous regions dominate the view, with the moving foreground instances only occupying a relatively small area [17]. However, the latter often exhibit higher stochasticity, complicating their prediction. Since Eq. 1 supervises all outputs uniformly, it cannot effectively discriminate the nuances of different regions as Fig. 4(b) shows. As a result, the model cannot efficiently learn to predict realistic dynamics in crucial regions.
@@ Appendix A Discussions
(1) As a generalizable predictive model, Vista could be utilized as a forward dynamics model [13, 25] to simulate short-term dynamics and assist long-horizon planning tasks like visual navigation [103]. (2) It is also intriguing to utilize Vista as an implicit driving policy, which is spontaneously acquired through future prediction [1, 24]. After synthesizing the video plan, we can convert the resultant image trajectory to executable actions by a non-causal inverse dynamics model [4, 68], which can be efficiently learned from much fewer data compared with the imitation learning pipeline [2, 9]. In autonomous driving, the inverse dynamics model could be implemented with visual odometry [72]. (3) In collaboration with the reward function, it is also worth investigating if Vista could facilitate model-based reinforcement learning by boosting the sampling efficiency in real-world scenarios [134].
@@
These two works have fundamental differences in control versatility and prediction fidelity. First of all, Vista is a generalizable world model that can be controlled by multi-modal action conditions. Although GenAD has also trained a trajectory-conditioned extension, its weights are fully finetuned on nuScenes and the generalization of its action control has never been validated. In contrast, Vista integrates versatile action controllability that can generalize to new scenarios in a zero-shot manner. Unlike GenAD that requires labeling OpenDV-YouTube with commands and texts, our collaborative training strategy skillfully averts this labor that may incur accumulated noises and conflicts [133]. In addition, Vista (10 Hz, 576$\times$1024) operates at much higher frame rate and resolution, considerably beyond the capability of GenAD (2 Hz, 256$\times$448) in both temporal and spatial axes. Different from GenAD, we also put forth several dedicated designs for high-fidelity prediction. We find that Vista, with a lower model complexity, achieves much better FID and FVD scores than GenAD (see Table 2).
@@
As one of the pioneering efforts, Vista still has a few limitations that call for future works. (1) Since Vista predicts futures at an exceptional spatiotemporal resolution, it is inevitable to be computationally expensive, particularly in downstream applications. Potential solutions may include faster sampling techniques [86, 143] and training-based distillations [88, 90, 98, 100, 107, 121]. (2) It is possible that the prediction may undergo an apparent degradation in long-horizon rollouts or drastic view shifts. Extra refinements on the prediction results [6, 48, 53, 96, 123, 141] could be helpful. Speculatively, applying our recipe to more scalable architecture [53, 94] is also promising to address this limitation. (3) Similar to other controllable video generation methods [125], the chance of failure still persists in our action controls, especially for ambiguous intentions such as commands and goal points as Fig. 8 reveals. Incorporating more datasets with action annotations [11, 109, 126] for collaborative training could be beneficial. Using compositional classifier-free guidance [8, 19, 31] to amplify the individual impact of action conditions may also help (at a cost of increased inference compute). (4) Although our training data is based on the largest public driving dataset [133], it is nowhere near the entirety of Internet driving data, thus leaving huge untapped potential to further expand the capabilities of Vista.
@@ B.1 World Models
Intelligent agents should be able to make efficacious decisions even under unseen situations [9, 56, 73, 110, 135, 146]. This requires fundamental knowledge of the world that generalizes to rare cases. As an internal manifestation of such knowledge, a world model predicts plausible futures of the world given potential actions [9, 39, 73, 93, 110, 134]. In principle, it not only predicts how the environment will unfold over time, but also deduces the underlying physical dynamics and agentic behaviors. Such properties can be useful for representation learning [34, 44, 85, 102, 128], model-based reinforcement learning [38, 39, 41, 42, 91, 93, 95], as well as model-predictive control [26, 29, 34, 40, 89, 101, 139]. Recent methods [36, 43, 77, 82, 83] also induce language-based world models from large language models, but are restricted in textual space and struggle with grounding on physics [25, 58].
Although world models have been extensively applied and made significant revolutions in simulated games [39, 41, 42] and indoor embodiment [69, 89, 117], such investigations for autonomous driving still lag behind [124, 140]. Different from other tasks, world modeling for autonomous driving poses unique challenges, which primarily arise from the large field of views with highly dynamic motions. Some practices imagine the world in the bird’s eye view (BEV) space [20, 30, 51, 52, 75, 78]. Recent practices model the world state as raw sensor observations such as point clouds [7, 65, 137, 140, 144] and images [53, 60, 66, 87, 99, 122, 124, 128, 142]. The latter category, namely visual world models, hold more promise for scaling up due to sensor flexibility and data accessibility. Nevertheless, existing methods are restricted to a particular dataset [60, 74, 87, 122, 124, 140, 142, 144] or simulator [7, 128], compromising their generalization ability to novel domains. Meanwhile, these efforts lack systematic designs for the driving domain and only model the world at relatively low frame rates and resolutions, which discards the fine-grained details and impairs their ability to express real-world behaviors. Moreover, most of them are restricted to a specific control modality [53, 60, 87, 122], which hinders the accommodation to prevailing planning algorithms [12, 14, 21, 55, 57, 63] and extension to more applications like decision-making [124] or user interaction [71]. Besides, existing methods seldom explore zero-shot action controllability across different datasets. The inferior generalization, fidelity and controllability collectively preclude existing driving world models from broadly facilitating the development of autonomous driving.
@@ C.2 Dataset
We utilize a rigorously filtered set of OpenDV-YouTube [133] for training, and incorporate nuScenes training set [10] during the action control learning phase. Concretely, we manually eliminate 15 hours of irrelevant content from OpenDV-YouTube, yielding approximately 1735 hours of unlabeled driving videos. Since nuScenes is heavily biased [80, 104, 124], we balance its samples based on command categories to foster the learning of rare actions. The video clips are sampled with 25 frames at 10 Hz. Although nuScenes [10] is logged at 12 Hz, we find no negative impact of treating them as 10 Hz videos. The model inputs are composed by cropping and resizing these clips to the target resolution. We define an action as a sequence comprising 25 frames. To categorize actions into commands, we follow the established conventions in planning [55, 57, 63] and define the command of ego-vehicle as "turn right" or "turn left" when its final displacement exceeds 2 meters in the orthogonal direction relative to its initial heading. To allow more precise categorization, we additionally introduce a "stop" command when the forward driving distance is less than 2 meters.
@@ D.2 Additional Ablation Studies
Table 4: Effect of action independence. Without losing generality, we choose the trajectory as a representative action for evaluation. The proposed constraint expedites the learning of actions.